OpenStudy (megannicole51):

Find the Taylor polynomials of degree n approximating ((4)/(5-5x)) for x near 0 For n=3, P3(X)= For n=5, P5(X)= For n=7, P7(X)= @agent0smith

OpenStudy (anonymous):

i would write it as \[\frac{4}{5}\frac{1}{1-x}\] and use the expansion of the geometric series , unless i am reading it wrong

OpenStudy (megannicole51):

how would you do that?

OpenStudy (anonymous):

\[\frac{1}{1-x}=1+x+x^2+x^3+...=\sum_{k=0}^{\infty}x^k\]

OpenStudy (anonymous):

a very frequent gimmick rather than taking successive derivatives which is usually a pain, is to arrange the thing you have to look like a geometric series somehow, or the derivative of one

OpenStudy (megannicole51):

so then what do i do with that series?

OpenStudy (anonymous):

multiply it by \(\frac{4}{5}\) then stop for whatever degree polynomial you want

OpenStudy (anonymous):

so for example \(P_3\) would be \[\frac{4}{5}(1+x+x^2+x^3)\]

OpenStudy (megannicole51):

and thats the answer?

OpenStudy (anonymous):

i think so, yes

OpenStudy (megannicole51):

yay it was right:)

OpenStudy (anonymous):

it is like all those problems you did where you had to sum a geometric series, only this time you are working backwards you have the closed form you write it in the open form

OpenStudy (anonymous):

whew, i am glad i didn't screw that one up

OpenStudy (megannicole51):

im definitely going to need some practice with these lol :/

OpenStudy (megannicole51):

thank you!

OpenStudy (anonymous):

yw

OpenStudy (megannicole51):

how would i do it with sin(3x)?

OpenStudy (anonymous):

memorize the power series for \(\sin(x)\) then replace \(x\) by \(3x\)

OpenStudy (anonymous):

do you know the power series for \(\sin(x)\) ?

OpenStudy (megannicole51):

no i don't:/

OpenStudy (anonymous):

ooooh that is a problem ok it is easy to remember first of all sine is odd, so the power series has only odd exponents

OpenStudy (anonymous):

and it alternates and the denominators are \(n!\) just like for \(e^x\)

OpenStudy (megannicole51):

okay:)

OpenStudy (anonymous):

\[x-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+...\]

OpenStudy (megannicole51):

oh wait doesnt it just alternate between 0,1,0,-1.....

OpenStudy (anonymous):

when i said alternate, i was ignoring the terms that are not there

OpenStudy (anonymous):

but yes, you can think of it that way if you like

OpenStudy (megannicole51):

i like urs better.....so then how do i do it when its 3x not just x? do i just add 3x to the numerator?

OpenStudy (anonymous):

oh i see what you're saying the derivatives evaluated as zero alternate between those numbers so yes, you know it and therefore you know the power series as well

OpenStudy (megannicole51):

oh no i dont know anything with this chapter lol it just sounded familiar! this is the worst chapter my professor has taught so far!

OpenStudy (anonymous):

\[sin(x)=x-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+...\] \[\sin(3x)=3x-\frac{(3x)^3}{3!}+\frac{(3x)^5}{5!}-\frac{(3x)^7}{7!}+...\]

OpenStudy (megannicole51):

yes that is exactly what i meant lol!

OpenStudy (anonymous):

yeah, then you are right i have to say this is tons easier than some of the stuff you have been doing

OpenStudy (megannicole51):

so its P1=3x?

OpenStudy (anonymous):

yeah

OpenStudy (megannicole51):

P3=3x-(((3x)^3)/(3!))+(((3x)^5)/(5!))

OpenStudy (megannicole51):

P5=3x-(((3x)^3)/(3!))+(((3x)^5)/(5!))-(((3x)^7)/(7!))+(((3x)^9)/(9!))

OpenStudy (anonymous):

i thought P3 meant the polynomial of degree 3

OpenStudy (anonymous):

not the one with three terms

OpenStudy (megannicole51):

oh shoot ur right.... so then how would i do it?

OpenStudy (anonymous):

degree 3 one is \[3x-\frac{(3x)^3}{3!}\]

OpenStudy (anonymous):

and degree 5 one is the one you wrote as the degree 3 one

OpenStudy (megannicole51):

how did u know that?

OpenStudy (megannicole51):

jk i see it now

OpenStudy (anonymous):

because \[3x-\frac{(3x)^3}{3!}\] is a polynomial... ok here i was going to explain it

OpenStudy (megannicole51):

thank you very much you were a lot of help:) so ahead and explain though!

OpenStudy (anonymous):

you might want to reduce the fraction and write is as \[3x-\frac{3x^3}{2}\]

OpenStudy (megannicole51):

okay:)

OpenStudy (anonymous):

you are welcome, and no i will not explain why \[3x-\frac{3x^3}{2}\] is a polynomial of degree 3 ! good luck with this, it is not that bad